### v1

# Nonequilibrium steady states in multi-bath quantum collision models

Abstract
Collision models provide a simple and versatile setting to capture the dynamics of open quantum systems. The standard approach to thermalisition in this setting involves an environment of independent and identically-prepared thermal qubits, interacting sequentially for a finite duration $\Delta t$ with the system. We compare this to a two-bath scenario in which collisional qubits are prepared in either their ground or excited states and the environment temperature is encoded in system-environment couplings. The system reaches the same thermal steady state for both settings as $\Delta t\to 0$, although even in this limit they describe fundamentally different physical processes, with the two-bath setup yielding a nonequilibrium state with finite heat currents. Non-Markovian dynamics arise when intra-environment interactions in either setting are introduced. Here, the system in the single-bath setup again reaches a steady state at the canonical temperature of the bath, but the nonequilibrium steady state of the two-bath setup tends to a different temperature due to the generation of strong system-environment and intra-environment correlations. The two-bath setting is particularly suited to studying quantum trajectories, which are well-defined also for the non-Markovian case. We showcase this with a trajectory analysis of the heat currents within a two-point measurement scheme. Our results provide insights into the dynamics and thermodynamics of thermalisation towards nonequilibrium steady states and the role of non-Markovian interactions.
I Introduction
A framework for understanding how a quantum system evolves when in contact with an environment is critical; both in terms of the practical development of quantum technologies and in addressing long-standing open questions regarding the dynamics of complex systems. Regarding the former noise characterisation in prototype quantum devices is crucial to overcome scalability issues Filippov et al. (2024), while the latter touches on major conceptual gaps in our understanding, e.g. determining the quantum-to-classical transition Zurek (2003) or how complex systems thermalise Deutsch (2018). It is not surprising then that a range of tools have been developed to simulate open quantum systems Purkayastha et al. (2021); Ferracin et al. (2024); Landi et al. (2022); Breuer and Petruccione (2002); Brenes et al. (2020). Within this plethora of techniques, collision models Ciccarello et al. (2022); Campbell and Vacchini (2021); Rau (1963); Scarani et al. (2002); Ziman et al. (2002) or repeated interaction schemes Strasberg et al. (2017), stand as a particularly versatile approach having been used to reconcile apparent thermodynamic inconsistencies present in other approaches De Chiara et al. (2018), explore how classicality emerges from an otherwise quantum dynamics Campbell et al. (2019); Ryan et al. (2022), and efficiently cool quantum systems Taranto et al. (2020); Langbehn et al. (2025).
The simplest formulation of a collision model involves imagining the environment as an infinite collection of identically prepared constituents, here after referred to as “auxiliary units” and labelled $A$, with which the system, $S$, sequentially interacts. Depending on the precise form of the $S$-$A$ interaction, and the initial states of the auxiliary units, this model accurately captures a range of Markovian dynamics. In particular, for so-called “thermal operations” it leads to thermalisation Scarani et al. (2002), while the model also more generally leads to homogenisation Ziman et al. (2002). However, varying the $S$-$A$ interaction can lead to a range of non-equilibrium steady states (NESS), whose characterisation has recently become the focus of several works Guarnieri et al. (2020); Heineken et al. (2021); Tian et al. (2021); Prositto et al. (2025); Palafox et al. (2022); Santiago-García et al. (2025). While assuming that all the auxiliary units are identically prepared is well justified, particularly in order to simplify simulation, this is clearly not a very realistic model. For instance, a more precise microscopic description of a thermal environment would assume that the system interacts with auxiliary units whose initial state is drawn in accordance with a distribution of energy that correspond to a given temperature. Several works have explored relaxing the assumption of identical auxiliary units and deterministic collisional events, leveraging scattering techniques to understand a wide class of open system dynamics Ehrich et al. (2020); Jacob et al. (2021); Tabanera-Bravo et al. (2023). The flexibility of collision models is clearly one of their inherent strengths. From the preceding discussion it is clear that collision models provide a useful approach to simulate how a system reaches the same, possibly thermal or equilibrium, steady state from different microscopic models.
This serves as the starting point for our analysis. We model a simple thermalisation process via collision models for two settings: the ‘typical’ situation where the auxiliary units are assumed to be all identical thermal states and a second setting where the system interacts simultaneously with two collisional baths with different initial energies, and these energies dictate the strength of the $S$-$A$ interaction. While these approaches are known to precisely recover the expected Markovian dynamics as described by the standard GKSL master equation, we show that they correspond to physically distinct dynamics, with the latter setting clearly driving the system to a non-equilibrium steady state. Through a simple modification of the basic collision model, we examine how the introduction of non-Markovianity to the dynamics significantly changes the steady state properties observed, with the establishment of strong system bath correlations playing a particularly important role. Finally, we explore the impact of introducing projective measurements on the auxiliary units using a two-point measurement scheme to recover the average heat current in the presence of non-Markovianity.
II Markovian Quantum Collision models
II.1 Single-bath homogenizing model (setting I)
We consider the typical collision model set-up which consists of a system of interest $S$, in an initial state $\rho_{0}$, interacting with an environment which is partitioned into individual constituents, the “auxiliary units”. These auxiliary units, labelled $A_{n}$, are each initialised in the state $\eta_{n}$. The time evolution is modeled in a stroboscopic manner by allowing the auxiliary units to ‘collide’ with the system for a fixed time $\Delta t$, such that a single interaction is given by the unitary operator (here and henceforth we set $\hbar=k_{B}=1$),
$U_{n}=e^{-i(H_{S}+H_{A_{n}}+H_{I})\Delta t}\;,$ (1)
where $H_{S}$ and $H_{A_{n}}$ are the free Hamiltonians of the system and auxiliary unit, respectively, and $H_{I}$ is the interaction Hamiltonian between $S$ and $A_{n}$. After $A_{n}$ has completed its interactions, we trace over its degrees of freedom. This procedure leads to the open system dynamics of $S$ described by the map,
$\rho_{n}=\Phi[\rho_{0}]=\mathcal{E}^{n}\circ\mathcal{E}^{n-1}\cdots\circ\mathcal{E}^{1}\left[\rho_{0}\right]\;,$ (2)
where $\circ$ indicates concatenation and,
$\mathcal{E}^{j}[\rho_{j-1}]=\text{Tr}_{j}[U_{j}(\rho_{j-1}\otimes\eta_{j})U^{{\dagger}}_{j}]\;,$ (3)
is the quantum channel describing the collision with the $j$-th auxiliary unit.
Collision models are a flexible framework for studying open system dynamics Ciccarello et al. (2022); Campbell and Vacchini (2021); indeed they can be used to simulate any time-continuous open system dynamics described by a Markovian quantum master equation Ziman et al. (2005). A simple setting of wide physical relevance is a two-level system with Hamiltonian, $H_{S}=\tfrac{\omega_{S}}{2}\sigma_{z}$, thermalising with a bosonic environment described by the master equation Landi et al. (2024),
$\dfrac{d\rho}{dt}=\mathcal{L}(\rho)=-i\left[H_{S},\rho\right]+\Gamma(\bar{N}+1)\mathcal{D}[\sigma^{-}]\rho+\Gamma\bar{N}\mathcal{D}[\sigma^{+}]\rho\;,$ (4)
where $\Gamma$ is the coupling strength and $\bar{N}\!=\!(e^{\beta\omega_{S}}-1)^{-1}$ is the Bose-Einstein distribution for a bath at (inverse) temperature $\beta\!=\!1/T$, as depicted in Fig. 1(a). The two dissipators $\mathcal{D}[\sigma^{-}],\mathcal{D}[\sigma^{+}]$, with $\sigma^{-},\sigma^{+}$ the Pauli lowering and raising operators, represent the emission and absorption of thermal photons, respectively. The steady state of Eq. (4) is given by the Gibbs state,
$\rho^{\text{th}}=\begin{pmatrix}\frac{1-g}{2}&0\\ 0&\frac{1+g}{2}\end{pmatrix}.$ (5)
where $g=(2\bar{N}+1)^{-1}$ Cherian et al. (2019).
The dynamics described by Eq. (4) can be stroboscopically simulated using a collision model Rau (1963); Scarani et al. (2002); Ciccarello et al. (2022), wherein the environment is viewed as an infinite collection of auxiliary units, each modeled as a qubit, identically prepared in a thermal state, i.e. $\eta_{n}=\rho^{\text{th}}$ given by Eq. (5), with the free evolution of each unit given by $H_{A_{n}}=\tfrac{\omega_{A_{n}}}{2}\sigma^{z}$. For simplicity, in the rest of this work we shall always assume the system and auxiliary units are on resonance, i.e. $\omega_{S}\!=\!\omega_{A_{n}}\!\equiv\!\omega$. This setting is illustrated in Fig. 1(b). In order to match the dynamics given by Eq. (4), we must fix the interaction Hamiltonian between the $S$ and $A_{n}$ to be,
$H_{I}=-\frac{J}{2}\left(\sigma^{x}_{S}\otimes\sigma^{x}_{A_{n}}+\sigma^{y}_{S}\otimes\sigma^{y}_{A_{n}}+\sigma^{z}_{S}\otimes\sigma^{z}_{A_{n}}\right)\;,$ (6)
with $J=\sqrt{{\frac{\Gamma(2\bar{N}+1)}{\Delta t}}}$. Scaling the interaction strength by $1/\sqrt{\Delta t}$ is necessary to ensure that Eq. (4) is recovered in the limit $\Delta t\to 0$. Throughout this work we will refer to this as ‘setting I’. Henceforth, we shall move to the interaction picture and neglect the free evolution of the system and auxiliary units. This can be done due to the fact that sum of the free Hamiltonians commute with the interaction Hamiltonian, i.e. $\left[H_{S}+H_{A_{n}},H_{I}\right]=0$  Lorenzo et al. (2017). This fact also means there is no switching work required to stoboscopically turn on and off the system-bath interaction De Chiara et al. (2018); Guarnieri et al. (2020). Under these conditions, the unitary in Eq. (1) corresponds to a partial-SWAP operation Scarani et al. (2002); McCloskey and Paternostro (2014); Campbell et al. (2018),
$U_{n}^{\text{I}}=\cos({J\Delta t})\mathbbm{1}-i\sin({J\Delta t})\mathcal{W}_{S,A_{n}}\;,$ (7)
where $\mathcal{W}$ is the SWAP operator,
$\mathcal{W}=\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\\ \end{pmatrix}.$ (8)
Figure 1: Schematic depiction of the approaches to model open quantum system dynamics considered in this work. (a) Typical open system dynamics where the system is interacts with a bath, generally assumed to be much larger than the system. Under certain assumptions this can be modeled using master equations. (b) Setting I: Collision models involve partitioning the environment into smaller constituent parts all prepared in the same initial state, $\eta_{i}$. The system interacts with a given environmental unit via some interaction with strength $J$. For interactions of the form Eq. (6) this collision model leads to homogenisation and the system reaches the steady state $\rho_{\infty}\!=\!\eta_{i}$. If $\eta_{i}$ are thermal states, this dynamics can be shown to be equivalent to the master equation, Eq. (4). (c) Setting II: Two bath collision model. The system interacts simulataneously with two independent collisional baths, where the constituents are initially in pure states |$\ket{0}$ and |$\ket{1}$, respectively, whose interaction strengths encode the effective temperature of the bath. In the limit of small interaction times, the system reaches the same steady state as setting I.
This set-up is a standard example of quantum homogenization, wherein the steady state of $S$ converges to the initial auxiliary state, i.e. $\rho_{\infty}\!=\!\eta_{i}$, for any choice of initial state for the auxiliary units $A_{n}$ Ziman et al. (2002). When $S$ sequentially interacts with $A_{n}$ via partial swap operations, its dynamics are described by a contractive completely positive map, guaranteeing convergence toward $\eta$ as the unique fixed point. This homogenization process is thermodynamically consistent and captures key features of thermalization. Importantly, in the limit of infinitely many collisions, the system not only reaches equilibrium but also asymptotically factorizes from the environment, resulting in a product state without residual system-environment correlations Cusumano et al. (2018).
II.2 Two-bath model (setting II)
An alternative approach to simulating the time-continuous dynamics of Eq. (4) involves allowing $S$ to interact with two collisional baths, one in which all the auxiliary units are in the |$\ket{0}$ state and one in which they are all in the |$\ket{1}$ state, where |$\ket{0}$ and |$\ket{1}$ corresponds to the ground and excited states of $A_{n}$, respectively. This setting is illustrated in Fig. 1(c). Throughout this work we will refer to this as ‘setting II’. The system interacts with a qubit from each bath simultaneously via an interaction Hamiltonian $H_{I}^{(i)}$ during a given collision, where we have introduced the superscript $i=0,1$ to label interactions with the baths |$\ket{0}$ and |$\ket{1}$, respectively. The interaction Hamiltonians take the same form as Eq. (6) but with different interaction strengths, $J^{(i)}$, which are given by
$J^{(0)}=\sqrt{\frac{\Gamma(\bar{N}+1)}{\Delta t}},\quad J^{(1)}=\sqrt{\frac{\Gamma\bar{N}}{\Delta t}}$ (9)
and encode the expected (effective) bath temperature. The unitary for a single step in this approach that enters into Eq. (1) is therefore (already in the interaction picture)
$U_{n}^{\text{II}}=e^{-i(H_{I}^{(0)}+H_{I}^{(1)})\Delta t}.$ (10)
In contrast to setting I, in this case the system will be driven to a non-equilibrium steady state (NESS), due to the simultaneous interaction with multiple inequivalent baths. This is because, irrespective of whether the system thermalizes or not, there will always be a constant flow of energy from the |$\ket{1}$ bath to the |$\ket{0}$ bath.
The unitary in Eq. (1) is also no longer given simply by a partial swap, and thus the system cannot homogenize. Similar to setting I however, we have $\left[H_{S}+H_{A_{n}},H_{I}^{(0)}+H_{I}^{(1)}\right]=0$ and so the free evolutions of the system and auxiliary units can again be neglected.
II.3 Steady States in the Markovian Limit
We start our analysis by noting that settings I and II both simulate the dynamics of a system in contact with a thermal bath. Indeed, the two settings are equivalent in the limit of $\Delta t\rightarrow 0$, where they both recover the dynamics and steady state that would be expected under the time continuous evolution given by Eq. (4). Furthermore, it is evident that they are equivalent in the zero temperature limit, $\beta\rightarrow\infty$. This is an obvious consequence of the fact that, in this limit, $J^{(1)}\!\to\!0$ and thus the two bath case reduces to the zero temperature limit of setting I. However, as alluded to previously, and as we will show explicitly below, setting II generally drives the system to a NESS.
The analytical form of the steady state in setting II is readily accessible, although too cumbersome to provide any direct insight. Nevertheless, it is diagonal in the system’s energy eigenbasis and, since we are considering a simple two-level system, it can therefore be assigned an effective temperature, $\beta_{e}$. This is formally defined as follows: as the state is diagonal, it can always be written in the form (5), with an effective parameter $g_{e}$. The effective temperature is then defined as
$\beta_{e}=\frac{1}{\omega}\log\left(\frac{1+g_{e}}{1-g_{e}}\right).$ (11)
As stated, the explicit formula of $g_{e}$ or $\beta_{e}$ is too cumbersome to write down. But, to illustrate the idea, in the case of very small $\Delta t$ we have
$g_{e}\simeq g+\frac{\Gamma\bar{N}(\bar{N}+1)}{6(2\bar{N}+1)^{2}}\Delta t\;.$ (12)
It therefore deviates from the equilibrium value $g=(2\bar{N}+1)^{-1}$ by a term that is proportional to $\Delta t$.
In Fig. 2 we show the deviation from the “canonical” steady state temperature, $\beta$, by calculating $\Delta\beta=\beta_{e}-\beta$. We clearly observe that in the limit of short collisions ($\Delta t\to 0$), the system is driven to the expected steady state, i.e. $\beta_{e}\!=\!\beta$, as predicted by Eq. (12). Increasing $\Delta t$, corresponding to longer interaction times during the collisions, leads to a substantial increase in $\Delta\beta$, with this deviation becoming more pronounced for increasing temperatures. Larger $\beta$ leads to larger deviations and we consistently find that $\Delta\beta\!>\!0$, indicating that the effective temperature that the system is driven to is cooler than expected. Considering each bath individually, it is clear that the $A_{n}^{(0)}$’s initialised in |$\ket{0}$ are at zero temperature, while the $A_{n}^{(1)}$’s in the |$\ket{1}$ can be considered as systems at negative infinite temperature. As the system is simulataneously in contact with both baths it follows that it must reach a NESS.
Maintaining the system in such a NESS requires heat currents to be flowing. Because the unitary is strictly energy preserving, there is no work associated with switching on and off the collisional unitaries De Chiara et al. (2018); Guarnieri et al. (2020). Hence all energy changes can be considered as heat exchanges. Focusing on the zero temperature bath in setting II, within the collision model setting we can readily calculate this via
$\Delta\mathcal{Q}^{(0)}(n)=\text{Tr}\left[H_{A_{n}^{(0)}}\left(\tilde{\rho}_{A_{n}^{(0)}}-\rho_{A_{n}^{(0)}}\right)\right]$ (13)
where $\tilde{\rho}_{A_{n}^{(0)}}$ is the state of the $A_{n}^{(0)}$ auxiliary unit after the system-environment collision has taken place. A similar expression holds for the $A_{n}^{(1)}$’s and, in fact, we find is exactly equal and opposite to Eq. (13), which is a direct consequence of strict energy conservation. The inset of Fig. 2 shows the steady state heat flux into the environment, i.e. $\frac{\Delta\mathcal{Q}}{\Delta t}\!\equiv\!\frac{\Delta\mathcal{Q}^{(0)}(n)}{\Delta t}$ for $n\!\rightarrow\!\infty$. We see that it is non-zero for all values of $\Delta t$. Smaller values of $\beta$ require larger heat fluxes to maintain the NESS. This is in line with expectation since in this regime the system is being driven to a state with a high effective temperature. It is therefore far from equilibrium from both of the respective collisional baths leading to substantial energy exchanges. Larger $\beta$ leads to a reduced heat flux since this regime tends to the zero temperature single bath situation captured by setting I. It is worth noting that the heat flux vanishes for all choices of $\beta$ in setting I Guarnieri et al. (2020).
Figure 2: Steady states in the Markovian limit. Main panel: Deviation of the effective temperature from the canonical temperature, $\Delta\beta$, with increasing collision time-step $\Delta t$ for different canonical temperatures $\beta\!=\!2$ (dashed, blue; lower temperature) and $\beta\!=\!0.5$ (solid, red; higher temperature). Inset: Non-equilibrium steady state heat current, $\Delta\mathcal{Q}/\Delta t$ for the initially low energy (cold) bath, $A_{n}^{(0)}$.
III Non-Markovian Quantum Collision Models
When using collision models to simulate Markovian dynamics, we make three assumptions regarding the constituents of the bath: (i) they do not interact with each other; (ii) they are initially uncorrelated; (iii) each constituent interacts with $S$ only once. Relaxing these assumptions allows us to extend collision models to simulate non-Markovian dynamics McCloskey and Paternostro (2014); Çakmak et al. (2017); Pezzutto et al. (2016); Zhang et al. (2019); Pleasance et al. (2025).
Here, we introduce non-Markovianity by breaking assumption (i) and introducing intra-bath collisions between consecutive units $A_{n}$ and $A_{n+1}$. We assume that this interaction occurs after the system has collided with $A_{n}$ but before it comes into contact with $A_{n+1}$. In all instances we fix this interaction to be
$U_{AA}=\text{cos}(\delta)\mathbbm{1}-i\text{sin}(\delta)\mathcal{W}_{A_{n},A_{n+1}}$ (14)
where $\mathcal{W}$ is as given in Eq. (8), and $0\!\leq\!\delta\!<\!\frac{\pi}{2}$. For the two settings considered previously, the unitary implementing the map is then
[…]
As before, the additional superscripts in setting II denote in which collisional bath the unitary acts. A sketch of these approaches is captured in Fig. 1(b) and (c) where the difference compared with the Markovian case is the inclusion of intra-environment collisions with some non-zero value of $\delta$.
The impact of intra-bath interactions is to correlate $S$ with $A_{n+1}$ before they have directly interacted with each other. This leads to a non-Markovian dynamics Breuer et al. (2016) with the parameter $\delta$ controlling the strength of the non-Markovianity McCloskey and Paternostro (2014). For $\delta\!=\!0$ the dynamics reverts to the Markovian case as discussed in Sec. II.3, while $\delta\!=\!\frac{\pi}{2}$ leads to an infinitely non-Markovian dynamics McCloskey and Paternostro (2014) since, for this interaction strength, Eq. (14) becomes a full SWAP and therefore the system is effectively interacting with the same qubit in the environment at every timestep and there is no loss of information.
A useful way to understand the introduction of the intra-bath interactions is through the notion of a memory-depth Campbell et al. (2018). For the Markovian collision models in Sec. II.3, each environmental constituent is involved in only a single collision and, thus, can be traced out after this interaction has taken place since it plays no role in the ensuing dynamics. Clearly, this is not the case for the non-Markovian setting as each $A_{n}$ is involved in two collisions, one with the system and a second with the incoming bath qubit, $A_{n+1}$. Introducing an additional full SWAP operation applied to $A_{n}$ and $A_{n+1}$ in Eqs. (15) allows to recover an effective Markovian dynamics Campbell et al. (2018); Kretschmer et al. (2016a). This is because the additional SWAP operation means that, in effect, the system interacts with the same bath unit, now considered a memory, during every collision. While this may seem like an unnecessary complication, it means that at the level of the system plus memory compound the dynamics is Markovian Campbell et al. (2018); Lorenzo et al. (2017); Kretschmer et al. (2016a). This Markovian embedding of non-Markovian dynamics allows to readily determine the system-memory steady state, from which the steady state of the system directly follows by tracing over the memory’s degrees of freedom.
Our interest is in characterising the steady states arising from the different microscopic models captured by Fig. 1. The ability to introduce non-Markovianity in via intra-environment interactions therefore adds an interesting additional dimension to our study, complementing recent works that focused on the Markovian limit Guarnieri et al. (2020); Prositto et al. (2025). To this end, we aim to assess the impact, if any, that degree of non-Markovianity plays in the resulting steady state properties. Several metrics have been proposed to quantify non-Markovianity Breuer et al. (2016); Addis et al. (2014). We shall focus on the BLP measure Breuer et al. (2009); Laine et al. (2010), which takes an information theoretic approach based on the behaviour of the trace distance between two states,
$D(\rho,\pi)=\frac{1}{2}||\rho-\pi||.$ (16)
Equation (16) is a quantifier of the distinguishability of two states, ranging from 1 for fully distinguishable states, to 0 for fully indistinguishable states. For a time continuous dynamics, the BLP measure is defined as
$N=\max_{\{\rho(0),\pi(0)\}}\int_{\Omega_{+}}\partial_{t}D(\rho(t),\pi(t))dt,$ (17)
where $\Omega_{+}=\bigcup_{i}(a_{i},b_{i})$ is the union of all the time intervals in the dynamics in which the trace distance between $\rho$ and $\pi$ increases, i.e. $\partial_{t}D(\rho(t),\pi(t))>0$. Here, $\rho(t)$ and $\pi(t)$ start from initially distinguishable (orthogonal) states. Equation (17) can be interpreted as summing over all the regions where the trace distance between a pair of system states increases during the dynamics, indicating an increase in their distinguishability and corresponding to an information backflow from the environment. The maximization is taken over all possible pairs of initial states. However, it has been shown that for qubit states, as is the focus of this work, the states that maximise Eq. (17) are always pure, orthogonal states Wißmann et al. (2012). For a collision model which simulates the open system dynamics via discrete time steps, we shall use a discretized version of Eq. (17) such that the non-Markovianity is quantified as McCloskey and Paternostro (2014); Laine et al. (2010),
$N=\max_{\{\rho_{0},\pi_{0}\}}\sum_{n}\left[D(\rho_{n},\pi_{n})-D(\rho_{n-1},\pi_{n-1})\right].$ (18)
III.1 Steady State Properties for Non-Markovian Collision Models
We begin by considering how the system approaches the steady state in each setting. For concreteness, we set $\beta\!=\!2$ which fixes initial auxiliary unit states for setting I and the interaction strengths, given in Eq. (9), for setting II. The most relevant parameters dictating the dynamics are then the $S$-$A_{n}$ interaction time $\Delta t$ and the strength of the inter-bath collisions, $\delta$. In Fig. 3 we examine the impact that these parameters have on the ensuing dynamics and the distance from the expected thermal state. We show the fidelity between the evolved state and Eq. (5),
$F=\left(\text{Tr}\left[\sqrt{\sqrt{\rho^{\text{th}}}\rho_{n}\sqrt{\rho^{\text{th}}}}\right]\right)^{2}.$ (19)
Figure 3: Dynamical approach to the steady state. Upper panel is for setting I and the lower panel is for setting II. We show the fidelity, ($1-F$), against $t\!=\!n\Delta t$, between the evolved system state and Eq. (5) fixing $\beta\!=\!2$. In both settings we consider three sets of parameters: Solid, red lines correspond to long system-environment collision durations, $\Delta t\!=\!0.01$, and strong intra-environment interactions $\delta\!=\!0.95\tfrac{\pi}{2}$; Dashed, blue curves correspond to long system-environment collision durations, $\Delta t\!=\!0.01$, and weaker intra-environment interactions, $\delta\!=\!0.8\tfrac{\pi}{2}$; Gray, dot-dashed curves correspond to short system-environment collision durations $\Delta t\!=\!0.001$ and strong intra-environment interactions, $\delta\!=\!0.95\tfrac{\pi}{2}$. The insets show the discretized non-Markovianity measure, Eq. (18), evaluated for $\beta\!=\!2$ as a function of intra-environment interactions strength $\delta$. In both main panels the system is initialized in the excited state $\rho_{0}=|1\rangle\langle 1|$. Other parameters used: $\omega=1,\Gamma=1.$
We remark that one could equally employ $\Delta\beta$ as considered in Sec. II.3 since throughout dynamics the system state remains diagonal and hence there exists a $\beta_{e}$. Indeed, we can explicitly relate the fidelity $F$ between $\rho^{\text{th}}$ for a given $\beta$ and $\rho_{n}$ to their temperature difference, $\Delta\beta$,
$F=\frac{\left(1+e^{\tfrac{1}{2}\omega(\Delta\beta+2\beta)}\right)^{2}}{\left(1+e^{\omega\beta}\right)\left(1+e^{\omega(\Delta\beta+\beta)}\right)}$ (20)
Considering first $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ for setting I (blue dashed line) we see that the system monotonically approaches the expected Gibbs state at temperature $\beta=2$. The dynamics appears effectively Markovian, despite the presence of the additional intra-bath interactions. Increasing the intra-bath interactions close to the maximal value is necessary before clear signatures of non-Markovianity emerge, as captured by the red solid curves in Fig. 3 where $\delta\!=0.95\tfrac{\pi}{2}\!$. However, these signatures depend on the value of $\Delta t$. Taking a smaller interaction time we find that, even in the presence of such extremely strong intra-bath interactions, the system’s dynamics still appears Markovian (gray dot-dashed line). Indeed, this can be confirmed by explicitly calculating the BLP measure, Eq. (18), shown in the inset. We see that for a given value of $\Delta t$, there is a critical value of $\delta$ below which the BLP measure is zero, and hence the dynamics appears Markovian. These dynamical differences notwithstanding, for all parameter choices in setting I the system is invariably driven to the expected thermal state. This can be readily confirmed by performing the Markovian embedding and calculating the steady state of system and memory Campbell et al. (2018). The nature of the system-bath interactions necessarily leads to homogenisation and, as will be relevant in later discussions, an asymptotic product state between system and bath units Cusumano et al. (2018).
Turning our attention to setting II, for the same parameters we find some notably different behaviors shown in the lower panel of Fig. 3. For $\delta\!=\!0.8\tfrac{\pi}{2}$ and $\Delta t\!=\!0.01$ (blue dashed line), the approach to the steady state is still monotonic, while for $\delta\!=\!0.95\tfrac{\pi}{2}$ (red line) the fidelity shows sizeable oscillations before settling to the steady state value. Thus, inline with setting I, strong intra-bath interactions are required to clearly signal the non-Markovian nature of the dynamics. We again find that there is a minimum value of $\delta$ below which the non-Markovian nature of the dynamics is not captured by revivals in the trace distance. We remark that while the BLP measure indicates that setting II is “more” non-Markovian, in the sense that smaller intra-bath collision strengths are necessary for revivals to be present, the overall trend is similar for the two settings and the amount of non-Markovianity as quantified by Eq. (18) is comparable. More interestingly, in the presence of such strong intra-bath interactions we see that the system’s steady state is substantially different from the expected thermal state. From the Sec. II.3 we know that even in the Markovian limit, setting II only thermalises to temperature $\beta$ precisely in the limit of $\Delta t\!\to\!0$, although the deviation is minor for sufficiently small, but strictly finite, values of $\Delta t$ (see Fig. 2). Here, however, we see that the non-Markovian dynamics leads to a significantly larger deviation from the expected thermal state.
We can conclude that settings I and II are essentially equivalent, in the sense that they yield the same thermal steady state, only in the Markovian case $\delta\!=\!0$, and in the limit of $\Delta t\!\to\!0$. Setting I, by construction, will drive the system to the Gibbs state at inverse temperature $\beta$ for any value of $\Delta t>0$ and $0\!\leq\!\delta\!<\!\tfrac{\pi}{2}$ Cusumano et al. (2018). By contrast, setting II drives the system to a NESS which is in general at a different temperature. The deviation of the NESS in setting II from the canonical thermal state becomes more significant as the degree of non-Markovianity is increased.
(a)                                                                   (b)
Refer to caption Refer to caption
(c)                                                                   (d)
Refer to caption Refer to caption
Figure 4: Panels (a) and (b): Difference in system temperature reached at the steady state, $\Delta\beta\!=\!{\beta_{e}-\beta}$ considering the impact of increasing the A-A interactions $\delta$ and increasing the S-A collision duration $\Delta t$. Panels (c) and (d): Associated steady state heat flux, $\Delta Q/\Delta t$, into $A_{n}^{(0)}$. Other parameters used: $\omega\!=\!1,\Gamma\!=\!1$.
In Fig. 4 we consider the impact of varying $\delta$ and $\Delta t$ on the resulting steady state in setting II, where we consider two values of temperature, $\beta\!=\!0.5$ and $2$. In panels (a) and (b) we show the difference $\Delta\beta$ between the effective temperature $\beta_{e}$ that the steady state reaches, and the canonical temperature $\beta$ appearing implicitly in Eq. (5), i.e. the Gibbs state that the system reaches in the limit of $\Delta t\!\to\!0$. We see a non-trivial interplay between the impact that the two parameters have on the resulting steady-state. For example, stronger non-Markovianity $\delta$ is required to see significant deviations from the expected thermal state when the collision time $\Delta t$ is small. We see that the regions of largest deviation correspond to values of $\Delta\beta\!<\!0$, i.e. the steady state is at a higher temperature than expected. This is markedly different from the Markovian case shown in Fig. 2 where $\Delta\beta\!>\!0$. Considering a larger value for $\beta$ itself, panel (b), we see a shrinking of the regions of most substantial deviation. These areas are generally pushed to higher values of $\delta$ and $\Delta t$, indicating that the non-equilibrium character of the steady state vanishes in the zero temperature limit. This is consistent with the fact that for $\beta\!\to\!\infty$, setting II reduces to the single bath case where homogenization is guaranteed. This variation in the effective temperature that the steady state reaches, corresponding to a transition from effective heating to cooling, implies that in the non-Markovian two bath setting the system can still reach the expected thermal state, shown by the white dashed lines in Fig. 4(a) and (b) indicating the parameter values where $\Delta\beta\!=\!0$. This demonstrates that while the non-Markovianity of the dynamics is a contributing factor for the system reaching a non-equilibrium steady state, there is no simple relationship between the degree of non-Markovianity and the effective temperature to which the system settles.
In the Markovian case $\delta=0$, Eq. (12) shows that in the limit $\Delta t\to 0$ the steady state of setting II is the canonical Gibbs state with $g_{e}\to g$ and hence $\beta_{e}\to\beta$. Interestingly, in the non-Markovian case, a different effective temperature $\beta_{e}\neq\beta$ can be reached even in the limit $\Delta t\to 0$. As can be appreciated visually from Fig. 4(a) and (b), the blue regions of large temperature deviation are pushed to larger and larger $\delta$ as $\Delta t$ becomes smaller. We therefore have a subtle order-of-limits issue, which can be analyzed by considering the effective temperature of the NESS as we send both $\Delta t\to 0$ and $\delta\to 1$ while keeping their ratio $r=\Delta t/(1-\delta)$ constant and finite. By recasting the problem in terms of variables $r$ and $\Delta t$ we find numerically that substantial deviations $\Delta\beta$ survive the $\Delta t\to 0$ limit for fixed values of $r>1$. However, we were not able to extract the non-Markovian analogue of Eq. (12) analytically. Our analysis shows that non-Markovian interactions have a significant effect on the dynamics in setting II, leading to a NESS with with a temperature different from the canonical temperature, even as $\Delta t\to 0$.
The energy-preserving nature of the interactions means that essentially the same heat current is observed due to all the collisions, i.e. the heat flux into the cooler bath is equal and opposite to the heat flux out of the excited state bath and, furthermore, the same magnitude of flux into and out of the auxiliary units is observed due to the intra-bath collisions. As a representative example, we show the heat flux into $A_{n}^{(0)}$ due to the system-bath interaction in Fig. 4(c) and (d). We see that the non-Markovian case exhibits large heat currents to maintain the NESS. As we will see in the following section, these currents allow for the establishment of strong correlations that elucidate why the system settles to a given effective temperature.
III.2 Steady State Correlations
We now consider the behavior of system-environment correlations. To do this we must employ a suitable measure and, for our purposes, we choose the negativity based on the postive partial transpose criterion. For a two qubit system, it is defined Peres (1996),
$\mathcal{N}_{2}=-2\max\left[0,\lambda_{neg}\right],$ (21)
where $\lambda_{neg}$ is the negative eigenvalue of the partially transposed density matrix. As we use the Markovian embedding to determine the steady state from the non-Markovian dynamics of setting II, our state consists of a three-qubit system made up of $S$ and the two “memory” qubits. We can extend Eq. (21) to quantify entanglement between different bi-partitions of the tripartite system. For three qubits $A,B$ and $C$ it follows that the entanglement between $A$ and the combined state of $B$ and $C$ is given by
$\mathcal{N}_{A(BC)}=-2\max\left[0,\sum_{j}\lambda_{neg,j}^{A(BC)}\right],$ (22)
where $\lambda_{neg,j}^{A(BC)}$ are the negative eigenvalues of the partially transposed density matrix of the bipartition $A$-$BC$. The genuine tripartite entanglement of the three-qubit state can then be calculated by taking a geometric average of the three possible bipartitions,
$\mathcal{N}_{3}=\left[\mathcal{N}_{A(BC)}\mathcal{N}_{B(AC)}\mathcal{N}_{C(AB)}\right]^{\frac{1}{3}}$
which provides an entanglement monotone. We remark that this corresponds to a sufficient, but not necessary, condition for entanglement.
In Fig. 5 we examine the entanglement established in the steady state of setting II in the full tripartite state as well as the various reduced state bipartitions. In all cases we find that there is a vanishingly small degree of entanglement present for sufficiently fast system-environment collisions and weak intra-bath interactions, i.e. small $\Delta t$ and $\delta$. Increasing these parameters allows for the establishment of entanglement in the tripartite system as shown in panel (a). The region of non-zero entanglement closely aligns with regions where setting II’s steady state significantly deviates from the expected thermal state. Considering the reduced state of the system and one of the baths, shown in panels (b) and (c), we again see a good agreement between regions where the system’s steady state is at a different effective temperature and the presence of bipartite entanglement between $S$ and $A_{n}^{(i)}$. For the reduced state between the two bath units, shown in panel (d), we see that entanglement is only established for large values of both $\Delta t$ and $\delta$. Interestingly, this region corresponds to the regions where $\Delta\beta$ is positive (see Fig. 4(d)), i.e. the establishment of cross-environmental entanglement aligns with the system being driven to cooler effective temperature.
(a)                                                                   (b)
Refer to caption Refer to caption
(c)                                                                   (d)
Refer to caption Refer to caption
Figure 5: Steady state entanglement arising from setting II. Panel (a): Genuine tripartite entanglement between the system and two “memory” units. Panels (b-d): Bipartite entanglement of the various reduced states. In all panels we fix $\beta\!=\!2,\omega=1,\Gamma=1$.
It is worth reminding that the above is in stark contrast to setting I’s steady state, which is driven to a factorised state between S and $A_{n}$ and therefore clearly there is no entanglement present in the steady state Cusumano et al. (2018). This can be intuitively understood due to the fact that for the considered interactions, setting I leads to homogenization and therefore in the steady state the system and all bath auxiliary units are all in identical states and, as the collisions correspond to a partial swap, they do not change the states. As demonstrated by Fig. 5, this is no longer the case in setting II, where the system evidently cannot homogenize and therefore each collisions is, generally, a strongly entangling operation.
We note that the entanglement in the tripartite NESS of system and memories is distinct from entanglement that is created during each collision of the memories with the incoming bath auxiliaries. Indeed, even in the Markovian case, once the system has reached the steady state, each subsequent collision in setting II will lead to the creation of genuine tripartite entanglement between the system and the bath units while leaving the reduced state of the system unchanged. This implies that the stronger deviation form the expected thermal state in the non-Markovian case is a consequence of setting II’s ability to create entanglement in the system memory compound in addition to the entanglement being continually created and destroyed in each subsequent collision with the baths. We therefore conclude that the presence of non-Markovianity enhances the non-equilibrium character of the state due to the map’s ability to create entanglement. However, we remark that as setting I has a comparable degree of non-Markovianity present in the dynamics (see Fig. 3), it follows that its presence is not solely responsible for the observed deviation from the expected thermal state and rather it is a more complex interplay between the constant energy exchanges, the creation and maintenance of strong quantum correlations, and the non-Markovian character of the dynamics.
IV Stochastic Heat Currents from Trajectories
Quantum collision models lend themselves naturally to the simulation of quantum trajectories when we consider performing measurements on the auxiliary units immediately after their interaction with the system Brun (2002). It is known that even for non-Markovian collision models, individual trajectories can be averaged to recover the unconditional dynamics Kretschmer et al. (2016b); Whalen (2019). In this section we will show how the unconditional heat current can be recovered from two-point measurements on the auxiliary qubits. The stochastic heat current after one complete collision is calculated by performing a projective measurement, with measurement operator $P_{j}$, in the energy eigen-basis of the auxiliary unit. This measurement is performed twice, first before the auxilliary unit has undergone any interaction with any other system, and again after it has completed all of its interactions. This procedure is commonly referred to as a two-point measurement (TPM) scheme. The post-measurement states of the auxiliary unit after the first and second measurements are given by,
$\rho_{A^{(i)}_{n}}^{\prime}=\frac{P_{j}\rho_{A_{n}^{(i)}}P_{j}}{\text{Tr}\left[P_{j}\rho_{A_{n}^{(i)}}\right]},\quad\rho_{A^{(.i)}_{n}}^{\prime\prime}=\frac{P_{j}\tilde{\rho}_{A_{n}^{(i)}}^{\prime}P_{j}}{\text{Tr}\left[P_{j}\tilde{\rho}_{A_{n}^{(i)}}^{\prime}\right]},$ (23)
where $\tilde{\rho}^{\prime}_{A_{n}^{(i)}}$ is the state of the auxiliary unit after it has undergone the first measurement and been involved in the the $A^{(i)}_{n-1}$-$A^{(i)}_{n}$,z $S$-$A^{(i)}_{n}$ and $A^{(i)}_{n}$-$A^{(i)}_{n+1}$ collisions. The stochastic heat current for the $n$-th time step is then
$\Delta\mathcal{Q}_{\text{stoc}}^{(i)}(n)=\text{Tr}\left[H_{A^{(i)}_{n}}\left(\rho^{\prime\prime}_{A_{n}^{(i)}}-\rho_{A^{(i)}_{n}}^{\prime}\right)\right].$ (24)
The unconditional heat current can be recovered from this stochastic current by averaging over many trajectories at each step $n$,
$\Delta\mathcal{Q}^{(i)}(n)\simeq\langle\Delta\mathcal{Q}^{(i)}(n)\rangle_{M}=\frac{1}{M}\sum_{k=1}^{M}\Delta\mathcal{Q}^{(i)}_{\text{stoc}}(n,k),$ (25)
where the averaging is done over $M$ trajectories. To give an illustrative example, in Fig. 6 we calculate the stochastic heat current for the zero temperature bath in setting II for two different values of $M$ and compare it to the unconditional heat current calculated using Eq. (13). It should be noted that since in this setting the zero temperature bath starts in the |$\ket{0}$ state, the first measurement in the TPM scheme has no effect and the TPM measurement scheme is equivalent to a typical continuous monitoring in this set-up. We see close agreement between the averaged stochastic heat and the unconditional value, with this agreement improving with increasing $M$.
Finally, it is interesting to note that certain heat fluxes in the non-Markovian setting, specifically those associated with the $S$-$A_{n}$ collisions, cannot be accessed using this type of TPM scheme. For example, in order to access the heat flux specific to the $S$-$A_{n}^{(0)}$ interaction a measurement must be performed on $A_{n}^{(0)}$ after the $S$-$A_{n}$ interaction, but before the $A_{n}^{(0)}$-$A_{n+1}^{(0)}$ interaction takes place. Clearly such a measurement can be done, however it will clearly affect the correlations shared between the $S$ and $A_{n+1}^{(0)}$ as well as the incoming state of $A_{n+1}^{(0)}$. This leads to a significant change in system dynamics and thus we will have a different stochastic head current compared to the unconditional dynamics. These “inaccessible currents” appear to be a unique feature of the non-Markovian setting.
Figure 6: Mean stochastic heat current into the zero temperature bath, $\langle\Delta\tilde{\mathcal{Q}}_{n}^{(0)}\rangle_{M}$ against $t\!=\!n\Delta t$ for $M=10,000$ (green, dashed line) and $M=100,000$ (black line) along with the corresponding unconditional heat current $\Delta\mathcal{Q}^{(0)}_{n}$ (magenta line). In all cases the system is initialized in the ground state $\rho_{0}=|0\rangle\langle 0|$. Parameters used: $\delta=0.95\frac{\pi}{2},\Gamma=1,\beta=1$.
V Conclusions
We have examined the steady state properties of a two-level quantum system undergoing an open dynamics simulated using quantum collision models. In the Markovian case, we considered a simple thermalisation process which can be effectively modelled via a standard GKSL master equation and considered two settings that precisely match this dynamics in the limit of vanishing collision time, $\Delta t\!\to\!0$.
As a benchmark we considered the typical situation, referred to as setting I, where the collisional bath consists of identical auxiliary units all prepared in the same thermal state. For the considered system-environment interactions, setting I is known to lead to thermalisation (and more generally homogenisation), leaving the system in equilibrium with the bath and, furthermore, the system and environment share no correlations in the steady state. By contrast, setting II involves two collisional baths, whose auxiliary qubits are prepared in either their ground or excited states, and the bath temperature is encoded through the system-bath couplings. As a main result we have demonstrated that despite the dynamics of the system being formally identical in the $\Delta t\!\to\!0$ limit, setting II necessarily drives the system to a nonequilibrium steady state (NESS), supported by non-zero steady state heat currents. By exploiting the flexibility of collision models, we examined the impact of non-Markovian dynamics on the resulting steady state, by introducing additional intra-environment interactions. Despite setting I and setting II exhibiting similar amounts of non-Markovianity, as quantified via the information back flow, we showed that the non-Markovian case of setting II results in the system being driven further from the expected equilibrium state. Although setting I always reaches the canonical temperature of the bath in the limit $\Delta t\to 0$, the non-equilibrium steady state of setting II can reach a different temperature. Through a careful analysis of the tripartite and bipartite entanglement, we have established that the non-equilibrium character of the steady state is principally rooted in the ability of setting II to create and maintain strong quantum correlations. We have also demonstrated how the average heat current can be recovered in setting II using a two-point measurement scheme, even in the presence of non-Markovianity.
Our results provide insight into the dynamics and thermodynamics of thermalisation and equilibration and highlight that seemingly equivalent dynamics at the level of the system of interest can correspond to fundamentally different physical processes. The versatility of the collision model framework allows them to be employed in the modeling of practical quantum devices and offers a promising avenue for future work. One possible direction is to model non-Markovian fermionic environments relevant to semiconductor quantum dot devices coupled to electronic circuits Goldhaber-Gordon et al. (1998); Shim et al. (2018); Barthelemy and Vandersypen (2013); Iftikhar et al. (2018); Pouse et al. (2023); Petropoulos et al. (2024), where we expect memory effects and strong system-bath correlations to play an important role. This builds on recent work showing that non-Markovian collision models are a powerful and systematic method for treating bosonic environments Lacroix et al. (2024). Another possibility would be to use the framework of non-Markovian collisional trajectories for measurement-based feedback control in non-Markovian environments, for example, to see the effects of specific intra-environment interactions and the resulting environmental memory on the ability to prepare a given target state Harwood et al. (2023).
Acknowledgements.
RMcE acknowledges support from Universities Ireland through the North/South Postgraduate Scholarship. RMcE and AKM acknowledge funding from Research Ireland through grant 21/RP2TF/10019. SC acknowledges support from the John Templeton Foundation Grant ID 62422.